OpenStudy (anonymous):

Solve -1/5x = 6 A. -30 B. -6/5 C. -1 D. 6 1/5

OpenStudy (ajprincess):

1.Multiply both sides by 5. 2.Divide both sides by -1

OpenStudy (phi):

from your previous problems you might have noticed: multiply both sides by the reciprocal of the number in front of the x you have -1/5 so multiply both sides by -5 (-1/5 flipped)

OpenStudy (phi):

you do this because -1/5 * -5 = 5/5 =1 and 1*x is x on the other hand (or should I say side) you get 6*-5 surely you can finish?

OpenStudy (anonymous):

I got 1

OpenStudy (anonymous):

I'm really confused now

OpenStudy (phi):

multiply BOTH sides by -5

OpenStudy (anonymous):

the ans is -1/30

OpenStudy (phi):

don't be confused start with \[ \frac{-1}{5} x = 6\] multiply both sides by the reciprocal of -1/5 which is -5 \[ -5 \cdot\frac{-1}{5} x = 6 \cdot -5\] can you finish?

OpenStudy (phi):

on the right side of the =, what is 6 times -5 ?